A heat lamp is used to treat a swollen leg. The lamp has an rms electric field of 3550 N/C. What is the maximum value of the electric field? What is the maximum intensity of the lamp? If the lamp if focused over a circular area of diameter 4.0 cm, what is the power delivered to the leg?
1
Expert's answer
2020-05-18T10:11:22-0400

Given data

rms value of electric field is "E_{rms}=3550 N\/C"

The maximum value of electric field is

"E_{max}=\sqrt{2}E_{rms}=\sqrt{2}(3550 N/C)=5020 N/C"

The intensity of electromagnetic wave is

"I=\frac{{E_{max}}^2}{2\mu_{0}c}=\frac{{5020}^2}{2(4\pi*10^{-7})(3*10^8)}=3.34*10^4 W/m^2"

The power delivered to the leg is

"P=IA=I\pi r^2=(3.34*10^4)(\pi(0.02 m)^2 =42 W"